OpenStudy (asad786):

On a $3000 loan, Bank A offers repayment terms of 6% simple interest for a 4-year term. Bank B offers 6% simple interest for a 5-year term. How much more interest will Bank B earn on the $3000 loan with the longer loan term?

OpenStudy (asad786):

@haleyelizabeth2017 @Mehek14

OpenStudy (photon336):

we need to apply the simple interest formula

OpenStudy (asad786):

hw?

OpenStudy (photon336):

\[I = P*R*T \] where I = how much the money grows with interest. P = principle starting amount of money R is the interest rate the bank offers. T is the amount of time in years.

OpenStudy (photon336):

now we've got two banks that are offering the same rate of 6% but for different amounts of time.

OpenStudy (photon336):

\[Bank~A = (\frac{ 6 }{ 100 })*(4~yr)(3,000) \] \[Bank~B = (\frac{ 6 }{ 100 })*(5~yr)(3,000) \] You would take the difference of the two to figure out how much more interest is gained by bank B.

OpenStudy (asad786):

i think answer is 180

OpenStudy (photon336):

yep you are right :)

OpenStudy (asad786):

thx